Part A A certain reaction has an activation energy of 67.0 kJ/mol and a frequency factor...

Part A

A certain reaction has an activation energy of 67.0 kJ/mol and a frequency factor of A1 = 1.40×1012 M−1s−1 . What is the rate constant, k, of this reaction at 28.0 ∘C ? Express your answer with the appropriate units. Indicate the multiplication of units explicitly either with a multiplication dot (asterisk) or a dash.

k =

Part B

An unknown reaction was observed, and the following data were collected: T (K) k (M−1⋅s−1) 352 109 426 185 Determine the activation energy for this reaction. Express your answer with the appropriate units.

Value
Answer #1

A)

Given that activation energy Ea = 67.0 kJ/mol = 67000 J/mol

frequency factor A = 1.40×1012 M−1s−1

temperature T = 28 C = 28 + 273 K = 301 K

Arrhenius equation is

k = A e-Ea/RT , k = rate constant

R = 8.314 J/K/mol

Then,

  k = A e-Ea/RT

   = (1.40×1012 M−1s−1) e[(- 67000 J/mol) /(8.314 J/K/mol) (301K)]

   = 3.3 M−1s−1

k = 3.3 M−1s−1

Therefore, rate constant k = 3.3 M−1s−1